the frequency of an organ pipe at 15degree c is 156hzat what tempearture will its frequency become 160hz

Grade:9

1 Answers

Khimraj
6 years ago
Frequency of organ is directly proportional to T3/2
So let at temperature T0C frequency of organ pipe become 160 then
(156/160) = (15 +273/T+273)3/2
then (288)/(T+273) = 0.983
So T +273 = 293
So T = 200C.
Hope it clears.